OpenStudy (anonymous):

Factor completely 200a^2+640ag+512g^2

OpenStudy (anonymous):

I get 8(5a+8g)^2 ?

OpenStudy (anonymous):

\[200a^2+640ag+512g^2=8(25a^2+80ag+64g^2)=8(5a+8g)^2\]Yeah, I think so, too.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No sweat.